Suppose an investment of $4,500 doubles in value every ten years. How much is the investment worth after 50 years? I got 45000 but i am not sure if i am correct.
Answer by Alan3354(69443)

Suppose an investment of $4,500 doubles in value every ten years. How much is the investment worth after 50 years?
-------------
50/10 = 5 doubling periods
2^5 = 32
$4500*32 = $144,000
--------------
Yrs Amt
10 9000
20 18000
30 36000
40 72000
50 144000